OpenStudy (anonymous):

What is the solution to the equation? 1/3x + 1/3 (2x - 15) = 3 1/2

Vocaloid (vocaloid):

start by distributing (1/3)(2x-15) = ?

OpenStudy (anonymous):

im confused

Vocaloid (vocaloid):

use the distributive property A(B-C) = A*B - A*C so (1/3)(2x-15) = ?

OpenStudy (anonymous):

so I would get x/32 -15... i think im doing this wrong

Vocaloid (vocaloid):

almost (1/3)*(2x) = ?

OpenStudy (anonymous):

wait so i think i got the answer... 41/2

OpenStudy (anonymous):

oh wait no umm... could you explain to me how to get the answer and step by step i have to leave soon but i need help so i know how to do this for any other test... this is the only question i need help with and my test will be over